OpenStudy (anonymous):

which is an equation in point-slope form for the given point and slope? point: (5,9); slope:2

OpenStudy (anonymous):

y = mx + b but m = 2 (slope), then from (5, 9), we get: 9 = 5*2 + b b = -1 y = 2x - 1

OpenStudy (anonymous):

in point slope form: (y - y0) = m(x - x0), where m is teh slope and (x0, y0) is a point in the line

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

No problem ;)